OpenStudy (anonymous):

Does anybody know how to do these? I am lost. Find the 3rd term in this sequence: Tn=2n --- 5

OpenStudy (anonymous):

put n=3 in the formula:)

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

T(3)=(2*3)/5=6/5

OpenStudy (anonymous):

but why do you put a 3?

OpenStudy (anonymous):

because its the 3rd term

OpenStudy (anonymous):

o ok now i get it thank you.

OpenStudy (anonymous):

good:)

OpenStudy (anonymous):

i have one more problem that i need help with. Can you help me please

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i want to see if i am doing it right.

OpenStudy (anonymous):

ok i am going to draw it

OpenStudy (anonymous):